# Understanding Codebase like a Professional! Human–AI Collaboration for Code Comprehension

### v2

@@ 8.1. Human-AI Co-understanding for Codebase
We conceptualize CodeMap as facilitating human–AI co-understanding of codebases (Figure 2), where AI acts not merely as an answer generator but as a collaborator that extracts, reformulates, and contextualizes information in ways aligned with human comprehension processes. Through this collaboration, the AI externalizes machine-level representations into human-readable structures, while humans iteratively steer the exploration through interpretation and inquiry. This loop turns comprehension into a shared activity rather than a one-way automation.
@@ 9. Threats to Validity
Internal validity. CodeMap ’s performance depends on the retrieval accuracy of its RAG component and the capabilities of the underlying LLM, including reasoning ability, context length, and susceptibility to hallucination. We mitigated these factors through iterative refinement and validated our approach using GPT-4.1, which demonstrated clear gains with improved model capabilities. Moreover, we involved 8 auditors for interviews, 15 participants (9 experienced, 6 novice) for validation, comparable to similar studies (e.g., 10 (Brachman et al., 2025), 12 (Xie et al., 2024)), but the sample size still limits representativeness. Larger and longer-term studies could further strengthen generalizability. Conclusion validity. We compared CodeMap only with ChatGPT and Understand. The effectiveness of other LLM-based tools, such as AI-powered IDEs (e.g., Cursor (Cursor, 2025)), remains uncertain. However, current AI IDEs mainly assist with context selection and lack CodeMap ’s hierarchical visualization and cognition support, limiting their effectiveness for deep code comprehension.
@@ 10. Related Work
LLM-powered Codebase Understanding. Recently, LLM-assisted snippet-level code understanding has attracted significant attention, including efforts to explain selected code snippets (Andryushchenko et al., 2024; Ross et al., 2023; Chen et al., 2023; Nam et al., 2024; Richards and Wessel, 2024) and related concepts (Nam et al., 2024; Andryushchenko et al., 2024). However, limited work have (Reefman, 2024; Berhanu Alebachew, 2025) explored the application of LLMs for codebase understanding, which is essential in many software engineering scenarios. Reefman et al. (Reefman, 2024) incorporate the LLM with static‑analysis context (call hierarchies, repository structure, etc.) to answer repository‑scale developer queries. However, because the system is implemented purely as a chatbot style interface, it lacks designs for human usage in real world. To address this gap, we adopted a human-centered perspective by examining how code auditors—who frequently onboard to new codebases—develop code comprehension strategies and how these insights can inform system design.
@@ 11. Conclusion
This study adopted a human-centered approach to explore how professional code auditors comprehend codebases and how their strategies can inform system design. We proposed four design opportunities for LLM-based systems and implemented CodeMap, a prototype that extracts, decomposes, and visualizes codebase information to support hierarchical understanding. User validation with experienced and novice developers showed that CodeMap enhances intuitiveness, ease of use, and usefulness, while reducing reliance on textual LLM responses. These findings demonstrate the potential of expert-informed, human–AI co-understanding systems to improve large-scale code comprehension and developer onboarding.
@@ 12. Data Availability
All data and code are available on our website (Authors, [n. d.]). This study has been approved by the Institutional Review Board (IRB) at the authors’ institution.

### v3

@@ 8.1. Human-AI Co-understanding for Codebase
We conceptualize CodeMap as facilitating human-AI co-understanding of codebases (Figure 2), where AI acts not merely as an answer generator but as a collaborator that extracts, reformulates, and contextualizes information in ways aligned with human comprehension processes. Through this collaboration, machine-level representations can be converted into human-readable structures, while humans iteratively steer the exploration through interpretation and inquiry. This loop turns comprehension into a shared activity rather than a one-way automation. CodeMap’s design shifts the interaction paradigm from standard prompting to UI-enhanced and context-based modes, as categorized in prior taxonomies of human-LLM interaction (Gao et al., 2024a).
@@ 9. Threats to Validity
Internal validity. CodeMap ’s performance depends on the retrieval accuracy of its RAG component and the capabilities of the underlying LLM, including reasoning ability, context length, and susceptibility to hallucination. We mitigated these factors through iterative refinement and validated our approach using GPT-4.1, which demonstrated clear gains with improved model capabilities. Moreover, we involved 8 auditors for interviews, 15 participants (9 experienced, 6 novice) for validation, comparable to similar studies (e.g., 10 (Brachman et al., 2025), 12 (Xie et al., 2024)), but the sample size still limits representativeness. Larger and longer-term studies could further strengthen generalizability. Conclusion validity. We compared CodeMap only with ChatGPT and Understand. The effectiveness of other LLM-based tools, such as AI-powered IDEs (e.g., Cursor (Cursor, 2025)), remains unexplored. However, current AI IDEs mainly assist with context selection and lack CodeMap’s hierarchical visualization and cognition support, limiting their effectiveness for deep code comprehension.
@@ 10. Related Work
LLM-powered Codebase Understanding. Recently, LLM-assisted snippet-level code understanding has attracted significant attention, including efforts to explain selected code snippets (Andryushchenko et al., 2024; Ross et al., 2023; Chen et al., 2023; Nam et al., 2024; Richards and Wessel, 2024) and related concepts (Nam et al., 2024; Andryushchenko et al., 2024). However, limited work has (Reefman, 2024; Alebachew, 2025) explored the application of LLMs for codebase understanding, which is essential in many software engineering scenarios. Reefman et al. (Reefman, 2024) incorporate the LLM with static‑analysis context (call hierarchies, repository structure, etc.) to answer repository‑scale developer queries. However, because the system is implemented purely as a chatbot style interface, it lacks designs for human usage in real world. To address this gap, we adopted a human-centered perspective by examining how code auditors–who frequently onboard to new codebases–develop code comprehension strategies and how these insights can inform system design.
@@ 11. Conclusion
This study adopted a human-centered approach to explore how professional code auditors comprehend codebases and how their strategies can inform LLM-powered Codebase Understanding System design. We proposed four design opportunities for LLM-based systems and operationalized them through CodeMap, a prototype that extracts, decomposes, and represents codebase information to support hierarchical understanding. User validation with experienced and novice developers showed that CodeMap enhances intuitiveness, ease of use, and usefulness, while reducing reliance on textual LLM responses. These findings demonstrate the potential of expert-informed, human–AI co-understanding systems to improve large-scale codebase comprehension and developer onboarding.
@@ 12. Data Availability
All data and code are available on our website. This study has been approved by the Institutional Review Board (IRB) at the authors’ institution.